OpenStudy (anonymous):

Use Gaussian elimination to find the complete solution to the system of equations, or state that none exists. 4x - y + 3z = 12, x + 4y + 6z = -32, 5x + 3y + 9z = 20, A.(8, -7, -2), B.(-8, -7, 9), C.none exist, D.(2, -7, -1).

OpenStudy (anonymous):

Please help need serious math help

OpenStudy (ybarrap):

$$ A=\begin{bmatrix} 4 &-1 &3 &| &12 \\ 1& 4 &6 &| &-32 \\ 5 &3 & 9 &| &20 \end{bmatrix} \\ \\ row(1)= -4*row(2)+row(1)\\ row(3)=-5+row(2)+row(3)\\ \rightarrow \begin{bmatrix} 0 &-17 &-21 &| &140 \\ 1& 4 &6 &| &-32 \\ 0 &-17 & -21 &| &180 \end{bmatrix} \\ row(1)=-row(3)+row(1)\\ \rightarrow \begin{bmatrix} 0 &0 &0 &| &-40 \\ 1& 4 &6 &| &-32 \\ 0 &-17 & -21 &| &180 \end{bmatrix} \\ \text{The 1st row means:}\\ 0x+0y+0z=-40\\ \text{There is no way that any x,y,z can produce -40. There are no solutions} $$